# Orthogonal Disentanglement with Projected Feature Alignment for Multimodal Emotion Recognition in Conversation

### v2

@@ Abstract
Multimodal Emotion Recognition in Conversation (MERC) significantly enhances emotion recognition performance by integrating complementary emotional cues from text, audio, and visual modalities. While existing methods commonly utilize techniques such as contrastive learning and cross-attention mechanisms to align cross-modal emotional semantics, they typically overlook modality-specific emotional nuances like micro-expressions, tone variations, and sarcastic language. To overcome these limitations, we propose Orthogonal Disentanglement with Projected Feature Alignment (OD-PFA), a novel framework designed explicitly to capture both shared semantics and modality-specific emotional cues. Our approach first decouples unimodal features into shared and modality-specific components. An orthogonal disentanglement strategy (OD) enforces effective separation between these components, aided by a reconstruction loss to maintain critical emotional information from each modality. Additionally, a projected feature alignment strategy (PFA) maps shared features across modalities into a common latent space and applies a cross-modal consistency alignment loss to enhance semantic coherence. Extensive evaluations on widely-used benchmark datasets, IEMOCAP and MELD, demonstrate effectiveness of our proposed OD-PFA multimodal emotion recognition tasks, as compared with the state-of-the-art approaches.
1 Introduction
@@
Multimodal Emotion Recognition in Conversation (MERC) enhances emotion recognition by integrating complementary emotion-related cues from multiple modalities, including text, audio, and visuals [jing2024dq, wang2024dynamic, ghosal-etal-2019-dialoguegcn, hu-etal-2021-mmgcn, fu2025himul]. This multimodal integration enables MERC to capture a wider spectrum of emotional expressions, resulting in significant performance gains over traditional unimodal approaches [jiang2024self, tu2022sentiment, zhang2023cross].
@@
To effectively integrate these cues from multiple modalities, most existing methods adopt contrastive learning [shou2024adversarial, li2024enhancing] and cross-attention [guo2024speaker, yang2023self, ai2024gcn] to capture their consistent semantics based on their similarity. For instance, MGCMA [shou2024adversarial] utilizes a distribution-based alignment module to exploit coarse-grained alignment representations among three modalities via contrastive learning. SACCMA[guo2024speaker] employs a cross-modal attention mechanism to maximize the consistent information from each modality. However, all above methods only focus on exploiting modality invariant information, which can be shared across all modalities, and ignore modality-specific emotion-related cues, such as micro-expressions, tone variations and sarcastic language.
@@
Recently, Joyful [li-etal-2023-joyful] employs a global contextual fusion module to extract global features shared across modalities, along with a specific fusion module to fuse modality-specific cues. However, neither its smoothing strategy nor reconstruction loss can explicitly align consistent semantics for global features. This limits its effectiveness in emotion recognition.
@@
To this end, we propose an Orthogonal Disentanglement and Projected Feature Alignment (OD-PFA) approach for the MERC task. Specifically, the unimodal features from each modality are first decoupled into shared and modality-specific components, enabling the model to simultaneously capture consistent semantics across all modalities and preserve the distinct cues unique to each modality. After this, we propose an orthogonal disentanglement (OD) strategy combined with a projected feature alignment (PFA) strategy to refine the extracted shared and specific features. Specifically, the OD strategy applies an orthogonal constraint to enhance the separation between shared and specific features, moreover, a unimodal feature reconstruction loss is incorporated to ensure that critical emotion-related information can be well preserved in both types of features of each modality. Meanwhile, the PFA strategy employs a feature projection operation to map the shared features of three modalities into a common latent space, and introduces a cross-modal consistency alignment loss to further improve the semantic consistency of these aligned shared features. Finally, the modality-specific features and the projected shared features are integrated to perform emotion recognition. Extensive experiments conducted on widely-used IEMOCAP [busso2008iemocap] and MELD [poria-etal-2019-meld] datasets demonstrate effectiveness of our proposed method, achieving comparable performance to the state-of-the-art approaches.
2 Proposed Method
@@
2.1 Multi-Modality Feature Extraction
2.1.1 Unimodal Feature Extraction
@@
2.1.2 Modality-Shared and Specific Feature Extraction
@@
2.2 Orthogonal Feature Disentanglement
@@
Inspired by [hazarika2020misa, li2023decoupled], we apply an orthogonal constraint to minimize the similarity between shared and specific representations, ensuring they capture distinct semantic aspects by enforcing a 90-degree angle between them. This constraint is applied both within the same modality and across different modalities. Therefore, the disentangle loss $\mathcal{L}_{\text{Dis}}$ can be calculated as,
$\mathcal{L}_{\text{Dis}}=\sum_{m\in\{l,v,a\}}(\mathbf{s}_{m,i}^{\top}\cdot\mathbf{p}_{m,i})^{2}+\sum_{\begin{subarray}{c}(m_{1},m_{2})\in\\ \{(l,a),(l,v),(a,v)\}\end{subarray}}(\mathbf{s}_{m_{1},i}^{\top}\cdot\mathbf{p}_{m_{2},i})^{2}$ (4)
@@
$\mathcal{L}_{\text{Rec}}=\frac{1}{3}\sum_{m\in\left\{l,a,v\right\}}\left\|\mathbf{x}_{m,i}-\hat{\mathbf{x}}_{m,i}\right\|^{2}_{2}$ (6)
2.3 Cross-Modal Feature Projection and Alignment
@@
2.3.1 Shared Space Feature Projection
@@
2.3.2 Cross-Modal Consistency Alignment
@@
Compared to pairwise alignment of shared features across all three modalities [hazarika2020misa, guo2024speaker, meng2024masked], our PFA strategy not only reduces alignment complexity, but also narrows modality gaps through projection operation, resulting in more efficient and effective alignment. Additionally, the cross-modal consistency alignment ensures that emotion related information is well-preserved and consistent across all modalities.
2.4 Emotion Recognition
For better recognition, we first enhance the projected shared feature of modality $m$ with contextual semantics by concatenating the projected shared features of all utterances within a conversation, and feeding the sequence into an $L$-layer Transformer (i.e., $\mathbb{T}_{m}(\cdot)$) with self-attention mechanism [vaswani2017attention] as follows,
@@ 2.4 Emotion Recognition
where the classifier $\mathbf{C}(\cdot)$ consists of a fully-connected layer and a subsequent 2-layer MLP with a ReLU, as in [shi2023multiemo].
@@
3 Experiments
3.1 Experimental Setup
@@
Dataset: We evaluate our method on two widely-used MERC benchmarks: IEMOCAP [busso2008iemocap] and MELD [poria-etal-2019-meld]. IEMOCAP contains $7,433$ utterances from $151$ dyadic dialogues across five sessions, annotated with six emotions: Happy, Sad, Angry, Excited, Frustrated, and Neutral. MELD consists of over $13,000$ utterances from over $1,400$ dialogues in the Friends TV series, labeled with seven emotions. We use the official dataset splits for all experiments.
@@
Evaluation Metrics: We adopt two widely used evaluation metrics for performance evaluation, including Accuracy (Acc) [yang2023self][guo2024speaker] and Weighted F1-score (w-F1) [hu2022mm] [li-etal-2023-joyful], following the state-of-the-art methods [meng2024masked, guo2024speaker].
3.2 Performance Comparison
@@
We compare OD-PFA with several state-of-the-art methods (i.e., DialogGCN [ghosal-etal-2019-dialoguegcn], MMGCN [wei2019mmgcn], MMDFN [hu2022mm] MGLRA[meng2024masked] and Joyful [li-etal-2023-joyful]) on both IEMOCAP and MELD datasets. Results are given in Table 1.
@@
Table 1: Performance comparison on IEMOCAP and MELD test dataset under the multimodal setting (text, audio and vision). Bold is the best, underlined is the second best. denotes the results from [hu2022mm] and others from original papers.
IEMOCAP
Methods Happy Sad Neutral Angry Excited Frustrated - Acc w-F1
DiaRNN [majumder2019dialoguernn] 32.20 80.26 57.89 62.82 73.87 59.76 - 63.52 62.89
MMGCN [wei2019mmgcn] 42.34 78.67 61.73 69.00 74.33 62.32 - 66.36 66.22
MMDFN [hu2022mm] 42.22 78.98 66.42 69.77 75.56 66.33 - 68.21 68.18
SCMM [yang2023self] 45.37 78.76 63.54 66.05 76.70 66.18 - - 67.53
Joyful [li-etal-2023-joyful] 60.94 84.42 68.24 69.95 73.54 67.55 - 70.55 71.03
SACCMA [guo2024speaker] 38.60 86.53 64.90 64.56 74.52 62.99 - - 67.10
MGLRA [meng2024masked] 63.50 81.50 71.50 61.10 76.30 67.80 - 71.30 70.10
OD-PFA(Ours) 66.90 80.09 71.58 64.46 79.93 67.75 - 72.09 72.34
MELD
@@
Methods Neutral Surprise Fear Sadness Joy Disgust Anger Acc w-F1
DiaRNN [majumder2019dialoguernn] 76.97 47.69 - 20.41 50.92 - 45.52 60.31 57.66
MMGCN [wei2019mmgcn] 76.33 48.15 - 26.74 53.02 - 46.09 60.42 58.31
MMDFN [hu2022mm] 77.76 50.69 - 22.93 54.78 - 47.82 62.49 59.46
SCMM [yang2023self] - - - - - - - - 59.44
Joyful [li-etal-2023-joyful] 76.80 51.91 - 41.78 56.89 - 50.71 62.53 61.77
SACCMA [guo2024speaker] - - - - - - - - 59.30
MGLRA [meng2024masked] 80.80 59.50 - 27.80 66.50 - 48.04 66.40 64.90
OD-PFA(Ours) 78.30 58.37 23.26 37.54 63.02 27.37 55.77 66.97 65.68
@@
It can be seen that our OD-PFA achieves the best performance in both Acc and w-F1 on IEMOCAP and MELD, especially excelling in recognizing Happy, Neutral, and Excited emotions on IEMOCAP, and Fear, Disgust, and Anger on MELD, demonstrating its strong effectiveness for MERC. Our OD-PFA performs better across both metrics than models like DialogRNN [majumder2019dialoguernn], SCMM [yang2023self], and SACCMA [guo2024speaker] that rely solely on cross-attention to capture consistent semantics. This underscores the importance of modeling both modality-shared consistency and modality-specific distinctions. In addition, OD-PFA surpasses Joyful [li-etal-2023-joyful], validating the effectiveness of our projected feature alignment strategy for aligning shared representations.
3.3 Ablation Study
@@
We analyze the effectiveness of our OD-PFA method by individually ablating the $L_{\text{Dis}}$ for disentangling shared and specific features, the $\mathcal{L}_{\text{PA}}$ for projected feature alignment, and the $\mathcal{L}_{\text{Cross}}$ for cross-modal consistency alignment. ‘w/o’ is applied to indicate without specified components. The results are presented in Table 2.
@@
Table 2: Validation of $L_{\text{Dis}}$, $\mathcal{L}{{}_{\text{PA}}}$ and $\mathcal{L}_{\text{Cross}}$ for MERC on both datasets.
@@
3.4 Analysis of Projected Shared Features
To further evaluate the effectiveness of the projection operation, we conduct an additional experiment to replace the projected features (i.e., $\mathbf{h}_{a,i},\mathbf{h}_{t,i},\mathbf{h}_{v,i}$) with original shared features (i.e., $\mathbf{s}_{a,i},\mathbf{s}_{t,i},\mathbf{s}_{v,i}$) in Section 2.4, named the ablated version as OD-PFA (w/o - Pro). The results are shown in Table 3.
Table 3: Validation of projected shared feature for MERC on both datasets.

### v3

@@ Abstract
Multimodal Emotion Recognition in Conversation (MERC) significantly enhances emotion recognition performance by integrating complementary emotional cues from text, audio, and visual modalities. While existing methods commonly utilize techniques such as contrastive learning and cross-attention mechanisms to align cross-modal emotional semantics, they typically overlook modality-specific emotional nuances like micro-expressions, tone variations, and sarcastic language. To overcome these limitations, we propose Orthogonal Disentanglement with Projected Feature Alignment (OD-PFA), a novel framework designed explicitly to capture both shared semantics and modality-specific emotional cues. Our approach first decouples unimodal features into shared and modality-specific components. An orthogonal disentanglement strategy (OD) enforces effective separation between these components, aided by a reconstruction loss to maintain critical emotional information from each modality. Additionally, a projected feature alignment strategy (PFA) maps shared features across modalities into a common latent space and applies a cross-modal consistency alignment loss to enhance semantic coherence. Extensive evaluations on widely-used benchmark datasets, IEMOCAP and MELD, demonstrate effectiveness of our proposed OD-PFA multimodal in emotion recognition tasks, as compared with the state-of-the-art approaches.
I Introduction
@@
Multimodal Emotion Recognition in Conversation (MERC) enhances emotion recognition by integrating complementary emotion-related cues from multiple modalities, including text, audio, and visuals [11, 23, 5, 9, 4]. This multimodal integration enables MERC to capture a wider spectrum of emotional expressions, resulting in significant performance gains over traditional unimodal approaches [10, 21, 26].
@@
To effectively integrate these cues from multiple modalities, most existing methods adopt contrastive learning [20, 14] and cross-attention [6, 25, 1] to capture their consistent semantics based on their similarity. For instance, MGCMA [20] utilizes a distribution-based alignment module to exploit coarse-grained alignment representations among three modalities via contrastive learning. SACCMA[6] employs a cross-modal attention mechanism to maximize the consistent information from each modality. However, all above methods only focus on exploiting modality invariant information, which can be shared across all modalities, and ignore modality-specific emotion-related cues, such as micro-expressions, tone variations and sarcastic language.
@@
Recently, Joyful [13] employs a global contextual fusion module to extract global features shared across modalities, along with a specific fusion module to fuse modality-specific cues. However, neither its smoothing strategy nor reconstruction loss can explicitly align consistent semantics for global features. This limits its effectiveness in emotion recognition.
@@
To this end, we propose an Orthogonal Disentanglement and Projected Feature Alignment (OD-PFA) approach for the MERC task. Specifically, the unimodal features from each modality are first decoupled into shared and modality-specific components, enabling the model to simultaneously capture consistent semantics across all modalities and preserve the distinct cues unique to each modality. After this, we propose an orthogonal disentanglement (OD) strategy combined with a projected feature alignment (PFA) strategy to refine the extracted shared and specific features. Specifically, the OD strategy applies an orthogonal constraint to enhance the separation between shared and specific features, moreover, a unimodal feature reconstruction loss is incorporated to ensure that critical emotion-related information can be well preserved in both types of features of each modality. Meanwhile, the PFA strategy employs a feature projection operation to map the shared features of three modalities into a common latent space, and introduces a cross-modal consistency alignment loss to further improve the semantic consistency of these aligned shared features. Finally, the modality-specific features and the projected shared features are integrated to perform emotion recognition. Extensive experiments conducted on widely-used IEMOCAP [2] and MELD [18] datasets demonstrate effectiveness of our proposed method, achieving comparable performance to the state-of-the-art approaches.
II Proposed Method
@@
II-A Multi-Modality Feature Extraction
II-A1 Unimodal Feature Extraction
@@
II-A2 Modality-Shared and Specific Feature Extraction
@@
II-B Orthogonal Feature Disentanglement
@@
Inspired by [7, 15], we apply an orthogonal constraint to minimize the similarity between shared and specific representations, ensuring they capture distinct semantic aspects by enforcing a 90-degree angle between them. This constraint is applied both within the same modality and across different modalities. Therefore, the disentangle loss $\mathcal{L}_{\text{Dis}}$ can be calculated as,
$\mathcal{L}_{\text{Dis}}=\sum_{m\in\{t,v,a\}}(\mathbf{s}_{m,i}^{\top}\cdot\mathbf{p}_{m,i})^{2}+\sum_{\begin{subarray}{c}(m_{1},m_{2})\in\\ \{(t,a),(t,v),(a,v)\}\end{subarray}}(\mathbf{s}_{m_{1},i}^{\top}\cdot\mathbf{p}_{m_{2},i})^{2}$ (4)
@@
$\mathcal{L}_{\text{Rec}}=\frac{1}{3}\sum_{m\in\left\{t,a,v\right\}}\left\|\mathbf{x}_{m,i}-\hat{\mathbf{x}}_{m,i}\right\|^{2}_{2}$ (6)
II-C Cross-Modal Feature Projection and Alignment
@@
II-C1 Shared Space Feature Projection
@@
II-C2 Cross-Modal Consistency Alignment
@@
Compared to pairwise alignment of shared features across all three modalities [7, 6, 17], our PFA strategy not only reduces alignment complexity, but also narrows modality gaps through projection operation, resulting in more efficient and effective alignment. Additionally, the cross-modal consistency alignment ensures that emotion related information is well-preserved and consistent across all modalities.
II-D Emotion Recognition
For better recognition, we first enhance the projected shared feature of modality $m$ with contextual semantics by concatenating the projected shared features of all utterances within a conversation, and feeding the sequence into an $L$-layer Transformer (i.e., $\mathbb{T}_{m}(\cdot)$) with self-attention mechanism [22] as follows,
@@ II-D Emotion Recognition
where the classifier $\mathbf{C}(\cdot)$ consists of a fully-connected layer and a subsequent 2-layer MLP with a ReLU, as in [19].
@@
III Experiments
III-A Experimental Setup
@@
Dataset: We evaluate our method on two widely-used MERC benchmarks: IEMOCAP [2] and MELD [18]. IEMOCAP contains $7,433$ utterances from $151$ dyadic dialogues across five sessions, annotated with six emotions: Happy, Sad, Angry, Excited, Frustrated, and Neutral. MELD consists of over $13,000$ utterances from over $1,400$ dialogues in the Friends TV series, labeled with seven emotions. We use the official dataset splits for all experiments.
@@
Evaluation Metrics: We adopt two widely used evaluation metrics for performance evaluation, including Accuracy (Acc) [25][6] and Weighted F1-score (w-F1) [8] [13], following the state-of-the-art methods [17, 6].
Settings: The batch size is set to 32 for both IEMOCAP and MELD. The model is trained for 100 epochs on IEMOCAP and 30 epochs on MELD. The Adam optimizer is adopted with $\beta_{1}=0.9$ and $\beta_{2}=0.99$.
III-B Performance Comparison
@@
We compare OD-PFA with several state-of-the-art methods (i.e., DialogGCN [5], MMGCN [24], MMDFN [8] MGLRA[17] and Joyful [13]) on both IEMOCAP and MELD datasets. Results are given in Table I.
@@
TABLE I: Performance comparison on IEMOCAP and MELD test dataset under the multimodal setting (text, audio and vision).
Methods IEMOCAP MELD
@@
Happy Sad Neutral Angry Excited Frustrated Acc w-F1 Neutral Surprise Fear Sadness Joy Disgust Anger Acc w-F1
DiaRNN[16] 32.20 80.26 57.89 62.82 73.87 59.76 63.52 62.89 76.97 47.69 - 20.41 50.92 - 45.52 60.31 57.66
MMGCN [24] 42.34 78.67 61.73 69.00 74.33 62.32 66.36 66.22 76.33 48.15 - 26.74 53.02 - 46.09 60.42 58.31
MMDFN [8] 42.22 78.98 66.42 69.77 75.56 66.33 68.21 68.18 77.76 50.69 - 22.93 54.78 - 47.82 62.49 59.46
SCMM [25] 45.37 78.76 63.54 66.05 76.70 66.18 - 67.53 - - - - - - - - 59.44
Joyful [13] 60.94 84.42 68.24 69.95 73.54 67.55 70.55 71.03 76.80 51.91 - 41.78 56.89 - 50.71 62.53 61.77
SACCMA [6] 38.60 86.53 64.90 64.56 74.52 62.99 - 67.10 - - - - - - - - 59.30
MGLRA [17] 63.50 81.50 71.50 61.10 76.30 67.80 71.30 70.10 80.80 59.50 - 27.80 66.50 - 48.04 66.40 64.90
OD-PFA (Ours) 66.90 80.09 71.58 64.46 79.93 67.75 72.09 72.34 78.30 58.37 23.26 37.54 63.02 27.37 55.77 66.97 65.68
@@
It can be seen that our OD-PFA achieves the best performance in both Acc and w-F1 on IEMOCAP and MELD, especially excelling in recognizing Happy, Neutral, and Excited emotions on IEMOCAP, and Fear, Disgust, and Anger on MELD, demonstrating its strong effectiveness for MERC. Our OD-PFA performs better across both metrics than models like DialogRNN [16], SCMM [25], and SACCMA [6] that rely solely on cross-attention to capture consistent semantics. This underscores the importance of modeling both modality-shared consistency and modality-specific distinctions. In addition, OD-PFA surpasses Joyful [13], validating the effectiveness of our projected feature alignment strategy for aligning shared representations.
III-C Ablation Study
@@
We analyze the effectiveness of our OD-PFA method by individually ablating the $L_{\text{Dis}}$ for disentangling shared and specific features, the $\mathcal{L}_{\text{PA}}$ for projected feature alignment, and the $\mathcal{L}_{\text{Cross}}$ for cross-modal consistency alignment. ‘w/o’ is applied to indicate without specified components. The results are presented in Table II.
@@
TABLE II: Validation of $L_{\text{Dis}}$, $\mathcal{L}{{}_{\text{PA}}}$ and $\mathcal{L}_{\text{Cross}}$ for MERC on both datasets.
@@
III-D Analysis of Projected Shared Features
To further evaluate the effectiveness of the projection operation, we conduct an additional experiment to replace the projected features (i.e., $\mathbf{h}_{a,i},\mathbf{h}_{t,i},\mathbf{h}_{v,i}$) with original shared features (i.e., $\mathbf{s}_{a,i},\mathbf{s}_{t,i},\mathbf{s}_{v,i}$) in Section II-D, named the ablated version as OD-PFA (w/o - Pro). The results are shown in Table III.
TABLE III: Validation of projected shared feature for MERC on both datasets.